The population data consists of 100 data values and has a mean value of 23.45 and standard deviation of 10. What is the standard deviation of this data set if this was a sample data?

a 0.99 x SQRT(10)

b.SQRT(99)

c.100/SQRT(99)

d.10*SQRT(99)
Answer #1

The sample standard deviation has a divisor of (n-1) while the poppulation standard deviation has a divisor of n.

Thus, sqrt(n)*sigma=sqrt(n-1)*s

Therefore, sigma=sqrt(99)*10/sqrt(100)

Therefore, sigma=sqrt(99)

Hence the standard deviation of this data is SQRT(99).

Please like the answer.